# Compositional disorder in a multicomponent non-reciprocal mixture: stability and patterns

### v2

@@ V Imaginary eigenvalues and steady state dynamics
In this section, we further characterize $\lambda_{0}$ for a finite number of components $N$ and relate the trends that we observe in the statistics of $\lambda_{0}$ to patterns in numerical solutions of the full nonlinear model in Eq. (II). $\lambda_{0}$ is real only when $N\to\infty$ a limit that represents a theoretical abstraction. In reality, the statistics of $\lambda_{0}$ exhibit strong finite size effects that are relevant to real experimental systems and dynamics observed in numerical simulations. The most important aspect of this is that $\lambda_{0}$ is imaginary with a significant probability - an important fact because many novel features of the multi-species NRCH model can be attributed to the emergence of complex eigenvalues in the linearised dynamics of a phase separating mixture.
@@
In Fig. 6(a) the mean value of $\lambda_{0}$ averaged over many realizations of the matrix $\mathbb{M}$ is shown for the exponential, discrete bimodal, and the identical distributions. The variance of $\lambda_{0}$, indicated by the error bar, is larger for the first two in comparison with the identical distribution. The important difference in the average value of $\lambda_{0}$ stems from the fact that in the absence of compositional disorder, $\mathbb{D}$ has only imaginary eigenvalues. For $\alpha\to 1$ we expect that $\mbox{Re}(\lambda_{0})=0$, meaning that all eigenvalues including $\lambda_{0}$ are now strictly complex such that the fraction of complex $\lambda_{0}$ approaches unity as seen in Fig. 6(b). In the other two cases, the diagonal disorder of $\mathbb{D}$ creates the possibility that $\lambda_{0}$ is real, meaning that the probability is capped by a limiting value as seen in Fig. 6(b). For the exponential, we expect that as $d\to 0$ meaning that the range of allowed values of $\bar{\phi}$ shrinks, the curve will bend upward and resemble the case with no compositional disorder.
Finally we show a few examples of the dynamics for a few chosen values of $\Theta$ below $\Theta_{\rm sp}$. We note that compositional disorder plays a role in the steady state dynamics as seen in Fig. 1(c) as seen in the difference between the kymographs representing the space-time dynamics for compositional disorder sampled from the bimodal distributions and the exponential in panels (a) and (b) respectively of Fig. 2. The same interaction matrix $\mathbb{M}$ has been chosen for the two panels (a) and (b) in Fig. 2. The parameters for the two choices of disorder are mention in the caption of the Fig. 2. The self-interaction $\Theta$ is lowered (see values indicated in the panels) to showcase the dependence of pattern on the difference $\Theta_{\rm sp}-\Theta$. We generically find spatiotemporal chaos in the steady state.
@@ VI Conclusions
To conclude, we have explored the role of compositional disorder in determining the onset of the spinodal instability and pattern formation in a random non-reciprocally interacting mixture. Using tools of random matrix theory, we have shown that the instability threshold goes down quadratically with the strength of non-reciprocity in a system with infinitely many species. The stability of a system with any compositional disorder is determined by the interplay of two parameters - the coefficient multiplying the square of the activity, and the threshold value of the effective temperature of the same system at zero activity. These parameters are determined solely by the distribution from which the disorder is drawn from. Through analytical arguments, we have motivated the dependence of this particular form and provided arguments for the relative stability of the active system compared to its passive counterpart. These predictions are validated through extensive numerical simulations of two types: (1) generating ensembles of the dynamical matrix to analyze the statistical properties of the most unstable eigenvalue, and (2) performing direct numerical simulations of the multi-species NRCH equation.
Previous work has shown that non-reciprocity plays a key role in tuning the system to the threshold of instability [36]. However, it is reasonable to assume that controlling the volume fractions of chemicals in a mixture—and thereby regulating phase separation—may be a more accessible experimental pathway. Our results, which demonstrate diverse dynamical behaviors arising from simple changes in the form of compositional disorder, are promising. This work represents a first step toward understanding the composition dependence of multi-component active mixtures. A detailed exploration of the resulting dynamical phases will be presented in future studies.
@@ Acknowledgements.
LP and SS acknowledge useful discussions and support from Ramin Golestanian and Navdeep Rana.
VII Appendix
@@
$G(z)=\int\mathrm{\leavevmode\nobreak\ d}z^{\prime}\frac{\rho\left(z^{\prime}\right)}{z^{\prime}-z}$ (27)
@@
$G(q)=\int\mathrm{\leavevmode\nobreak\ d}z^{\prime}\rho\left(z^{\prime}\right)\left(z^{\prime}-q\right)^{-1}$ (28)
@@
$g+\beta\mathrm{j}=\int\mathrm{\leavevmode\nobreak\ d}\bar{\phi}P(\bar{\phi})\frac{(\bar{\phi}-{z^{*}}-(1-2\alpha^{2})g^{*})+\beta\mathrm{j}}{|\bar{\phi}-z-(1-2\alpha^{2})g|^{2}+|\beta|^{2}},$ (29)
where both sides of this equation are quaternionic. Writing the complex number $g$ as $g=\text{Re}(g)+\text{Im}(g)$ distinguishing the real and the imaginary parts, and $z=x+yi$, we can distinguish the real and imaginary parts of Eq. (29) to obtain the coupled equations for $z$ and $\mbox{Re}(g)$, see Eqs. (11) and (12), in the main text. In arriving at Eqs. (11) and (12) we have further assumed that the diagonal entries of $\mathbb{C}$, and thus the values $\bar{\phi}$ assumes are real.
VII.1 Numerical schemes and parameter choices
In this section, we will describe the methods used to solve the $N$ nonlinearly coupled PDEs to verify the results of the linear stability analysis and to garner some idea of the rich pattern forming ability of these systems of equations.
The simulations shown in Fig. 1 and used in Fig. 4 have been performed using a pseudo-spectral method in a one-dimensional system on a domain of length L discretized with nx grid points to obtain a stable solution of the nonlinear partial differential equations for the N-species equations of motion. For time discretizing, we use the second-order Exponential Time Differencing (ETD2) scheme [70] with a fixed time step $\Delta t$, a small enough time step to ensure reaching a steady state.
For the results in the Fig. (2), the simulation was carried out for $N=100$ components. The system size is chosen to be $L=4096$ and $nx=4096$ in the simulations. The time step for simulations reported in this figure is $\Delta t=2.5\times 10^{-3}$.
The results of the Fig. (6), is for a simulation of $N=100$ components with the system size $L=512$ and $nx=512$ in the simulations. The time step for simulations reported in this figure is $\Delta t=5\times 10^{-3}$. For both results in Fig (2) and (6) the surface tension parameter $K=0.1$.
In all simulations for the discrete bimodal distribution in we chose $d_{1}=0.2$ and $d_{2}=1$. Accordingly, for the exponential distribution that was used in the simulation, we chose $d=0.6$.
In Fig. (6), for the discrete bimodal distribution, from the linear stability analysis we can find $\Theta_{sp}=0.06$ and for the exponential distribution is $\Theta_{sp}=-0.19$.
@@ VII.2 Discrete Bimodal distribution
As a first example, we sample the diagonal entries of $\mathbb{C}$ from the bimodal distribution.
@@
We assume $d_{1}<d_{2}$ and $d_{1}>0$. Eqs. 11 and 12 assume simple forms for this choice, such that $\lambda_{0}$ can be found analytically setting $y=0$, and using the re-defined variable $\tilde{x}$ we get
@@
The smallest eigenvalue is then obtained as $\lambda_{0}=\tilde{x}-d_{2}-(1-2\alpha^{2})\sigma^{2}\text{Re}(g)$. However, it is worth mentioning that the boundary of support plotted in the 3 is for $\alpha=\sqrt{1/2}$ for which the boundary follows a simpler equation

### v3

@@ V Imaginary eigenvalues and steady state patterns
In this section, we further characterize $\lambda_{0}$ for a finite number of components NN and relate the trends that we observe in the statistics of $\lambda_{0}$ to patterns in numerical solutions of the full nonlinear model in Eq. (II). $\lambda_{0}$ is real only when $N\to\infty$, a limit that represents a theoretical abstraction. In reality, the statistics of $\lambda_{0}$ exhibit strong finite-size effects that are relevant to real experimental systems and to the dynamics observed in numerical simulations. The most important aspect of this is that $\lambda_{0}$ is imaginary with a significant probability—an important fact because the novel features of the multi-species NRCH model can be attributed to the emergence of complex eigenvalues in the linearised dynamics of a phase-separating mixture [33].
@@
In Fig. 5(a), the mean value of $\lambda_{0}$ (both the real and imaginary parts), averaged over 300 realisations of the matrix $\mathbb{M}$, is shown for the exponential, the discrete bimodal, and the identical distributions. In panel (b) of the same figure, the probability of $\lambda_{0}$ being complex is plotted for the same ensemble. A few comments are in order. First, the variance of $\lambda_{0}$, indicated by the error bars, shows that finite-size effects are larger for the first two cases in comparison with the identical distribution. Secondly, in the absence of compositional disorder, $\mathbb{D}$ has almost all imaginary eigenvalues for $\alpha\to 1$. The fraction of complex $\lambda_{0}$ approaches unity as NN approaches infinity, a relevant regime relevant to bio-condensates, as seen in Fig. 5(b). In the other two cases, the varying diagonal entries of $\mathbb{D}$ create the possibility that $\lambda_{0}$ is real, meaning that the probability that $\lambda_{0}$ is imaginary is capped by a limiting value, as seen in Fig. 5(b). For the exponential distribution, we expect that as $d\to 0$, or as the range of allowed values of $\bar{\phi}$ shrinks, the curve will bend upward and resemble the case with no compositional disorder.
Figure 7: Steady state dynamics for mixture (composition drawn from an exponential distribution): The left panels illustrate the dynamics in the steady state with the diagonal disorder from an exponential distribution. Each kymograph shows the evolution of the scalar field for one species among NN of them. At $\Theta$ close to $\Theta_{\rm sp}$, the dynamics are chaotic, resembling chaotic condensates. As $\Theta$ is decreased, the chaotic nature of the condensates increases. The dynamics is reflected in the histograms in panel (b) where peaks in the distribution of the values assumed by the fields changes to smooth curves with two well-defined peaks.
We now show a few examples of the dynamics in the steady state for values of $\Theta$ chosen to be below $\Theta_{\rm sp}$. In Figs. 6, Figs. 7 illustrate the dynamics in the steady state for the bimodal distribution, and the exponential distribution as $\Theta$ is tuned below $\Theta_{\rm sp}$. The same interaction matrix $\mathbb{M}$ has been chosen for the examples in panel (a) of both Figs. 6 and 7, the other relevant parameters are mentioned in the captions.
The effective temperature $\Theta$ is lowered (see values indicated in the panels) to showcase the dependence of pattern on the difference $\Theta_{\rm sp}-\Theta$. The kymographs representing the space-time dynamics for a randomly selected species in panel (a) of Fig. 6 shows that just below $\Theta_{rmsp}$ condensates appear. The term condensates refers to spatial domains of enhanced or diminished density that are observed that are long lived (being stationary in the steady state). For $\Theta=-0.1$, the system seems to have arrested, a consequence of the dimensionality, and deterministic evolution of the system. On lowering the value of $\Theta$, the condensates become dynamic, notice that they shift and fluctuate in time. On increasing $\Theta$ further, we generically find spatiotemporal chaos in the steady state.
To ensure that the system has reached steady state we calculate the distribution of the values assumed by the fields in the steady state at different points in the lattice and average the data temporally as well to obtain the histograms in panel (b) of both Figs 6 and 7. The formation of two sharp peaks would signal binary phase separation, while smooth peaks would correlate with spatiotemporally chaotic dynamics. Notice the correlation between panels (a) and (b), stationary condensates are associated with multiple well-defined peaks in the histogram, while chaotic ones are associated with smooth and wide peaks. We note that the solid lines in the histogram plots are not a parametric fit assuming any functional form but a result of data-driven smoothing to help in distinguishing the peaks. Interestingly, we find that the condensates are likely to represent multiphase coexistence, as seen in the three or more prominent peaks are observed for $\Theta_{\rm sp}$ just below $\Theta_{\rm sp}$.
Finally, A comparison between the dynamics in Figs. 6 and Fig. 7 shows that similar trends appear in both cases although the transition to a chaotic condensates occurs at a larger value of $\Theta$ for the exponential distribution. Although we have reported a few examples here, further simulations have shown that the trends are quite generic when other realisations of the interaction matrix $\mathbb{M}$ are considered.
@@ VI Conclusions
To conclude, we have explored the role of compositional disorder in determining the onset of the spinodal instability and pattern formation in a random nonreciprocally interacting mixture. The spinodal threshold (the temperature at which the system transitions from a homogeneous to a patterned state) in a system with varying mean density of each component decreases quadratically with the strength of the nonreciprocal parameter. The stability is thus governed by the interplay of two parameters. The first is the coefficient multiplying the square of the nonreciprocal activity, and the second is the transition point at zero nonreciprocity. Both parameters are determined solely by the distribution from which the disorder is drawn. Using analytical arguments based on random matrix theory, we have motivated this functional dependence and provided arguments for the relative stability of the active system compared to its passive counterpart.
These predictions are validated through extensive numerical simulations probing whether the onset of pattern formation follows the theoretical expectations. A second set of simulations was carried out to explore the emergent features of the steady-state dynamics. We find that condensates (static domains of enhanced density) appear close to the threshold of the spinodal instability and exhibit signatures of multi-phase coexistence, meaning that the system separates into more than two dominant phases, as evidenced by multiple peaks in the distribution of density values in the steady state. On lowering the effective temperature the condensates become chaotic, a feature common to both types of compositional explored in the paper.
In passive multicomponent systems, tuning the mean density of each species generally leads to complex phase behaviour, with changes in composition providing access to new critical points [22]. It is therefore striking that relatively simple statements can be made about the stability of the nonreciprocal system (arguably a much more complex system) when one considers an ensemble of systems with interaction strengths drawn from random distributions. The generality of the results in this realistic setting, with many components each having a variable average density, suggests that tuning the level of nonreciprocity can be an effective way to control the nature of condensates in biological systems [77]. However, it is reasonable to expect that controlling the volume fractions of chemical species in a mixture, and thereby regulating phase separation, may provide a more accessible experimental pathway [15].
Our work represents a first step toward understanding the composition dependence of multicomponent active mixtures, addressing the important aspect that different species are mixed in asymmetric proportions in realistic experimental setups and real biomolecular condensates. The detailed phase behaviour via a construction of binodals, and an implementation of Maxwell construction will be left for future work [52]. A detailed exploration of the resulting dynamical phases in two and three dimensions, where novel interfacial effects are expected upon changing the concentrations of the components [52], will be presented in future studies.
@@ Acknowledgements.
LP and $S$ acknowledge useful discussions and support from Ramin Golestanian and Navdeep Rana.
VII Appendix I
@@
$G(z)=\int\mathrm{\penalty 10000\ d}z^{\prime}\frac{\rho\left(z^{\prime}\right)}{z^{\prime}-z}$ (30)
@@
$G(q)=\int\mathrm{\penalty 10000\ d}z^{\prime}\rho\left(z^{\prime}\right)\left(z^{\prime}-q\right)^{-1}$ (31)
@@
$g+\beta\mathrm{j}=\int\mathrm{\penalty 10000\ d}\bar{\phi}P(\bar{\phi})\frac{(\bar{\phi}-{z^{*}}-(1-2\alpha^{2})g^{*})+\beta\mathrm{j}}{|\bar{\phi}-z-(1-2\alpha^{2})g|^{2}+|\beta|^{2}},$ (32)
where both sides of this equation are quaternionic. Writing the complex number $g$ as $g=\text{Re}(g)+\text{Im}(g)$ distinguishing the real and the imaginary parts, and $z=x+yi$, we can distinguish the real and imaginary parts of Eq. (32) to obtain the coupled equations for $z$ and $\mbox{Re}(g)$, see Eqs. (18) and (17), in the main text. In arriving at Eqs. (18) and (17) we have further assumed that the diagonal entries of $\mathbb{C}$, and thus the values $\bar{\phi}$ assume are real.
VII.1 Identical distribution
For $P=\delta(\bar{\phi}-d)$, Eqs. (18) and (17) reduce to
$\frac{\sigma^{2}/4}{(\tilde{x}-d)^{2}+\tilde{y}^{2}}=1,$
$\mbox{Re}(g)=\frac{d-\tilde{x}}{(\tilde{x}-d)^{2}+\tilde{y}^{2}}.$ (33)
which yields $\tilde{x}=d-\sqrt{\sigma^{2}/4-\tilde{y}^{2}}$, such that for $\tilde{y}=0$, is $\tilde{x}=d-\sigma/{2}$. $\mbox{Re}(g)=2/\sigma$. Using the definition of $\tilde{x}$. we have
$\lambda_{0}=d-\sigma(1-\alpha^{2}).$ (34)
@@ VII.2 Discrete Bimodal distribution
As the second example, we sample the diagonal entries of $\mathbb{C}$ from the bimodal distribution.
@@
We assume $d_{1}<d_{2}$ and $d_{1}>0$. Eqs. 18 and 17 assume simple forms for this choice, such that $\lambda_{0}$ can be found analytically setting $y=0$, and using the re-defined variable $\tilde{x}$ we get
@@
The smallest eigenvalue is then obtained as $\lambda_{0}=\tilde{x}-d_{2}-(1-2\alpha^{2})\sigma^{2}\text{Re}(g)$. However, it is worth mentioning that the boundary of support plotted in the 3 is for $\alpha=1/\sqrt{2}$ for which the boundary follows a simpler equation
@@
VIII Appendix II: numerical schemes and parameter choices
In this section, we will describe the methods used to solve the NN nonlinearly coupled PDEs to verify the results of the linear stability analysis and to garner some idea of the rich pattern forming ability of these systems of equations. The simulations shown in Fig. 1 and used to obtain the results in Fig. 2 have been performed using a pseudo-spectral method in a one-dimensional system on a domain of length $L$ discretized with nxnx grid points to obtain a stable solution of the nonlinear partial differential equations for the N-species equations of motion. For time discretizing, we use the second-order Exponential Time Differencing (ETD2) scheme [81] with a fixed time step $\Delta t$, a small enough time step to ensure reaching a steady state.
For the results in the Fig. 2, the simulation was carried out for $N=100$ components. The system size is chosen to be $L=4096$ and $nx=4096$ in the simulations. The time step for simulations reported in this figure is $\Delta t=2.5\times 10^{-3}$.
The results of the Figs. 6 and 7, are from a simulation of $N=100$ components with the system size $L=512$ and $nx=512$ in the simulations. The time step for simulations reported in this figure is $\Delta t=5\times 10^{-3}$. For both results in Fig. (2) and (6) the surface tension parameter $K=0.1$. In all simulations for the discrete bimodal distribution in we chose $d_{1}=0.2$ and $d_{2}=1$. Accordingly, for the exponential distribution that was used in the simulation, we chose $d=0.6$. In Fig. (6), for the discrete bimodal distribution, from the linear stability analysis, we can find $\Theta_{sp}=0.06$ and for the exponential distribution is $\Theta_{sp}=-0.19$.